# Abdoul Kader Kaboré

### v3

@@ 3.3 Large Language Models
Furthermore, all models are set to a standard temperature of 0.7. Our selection criteria prioritized models that excel in code generation and possess accessible APIs, making them suitable for both academic research and industrial applications (Wang et al., 2024; Fan et al., 2023).
@@ 3.4 Prompt Design
This study evaluates five prompt techniques: Zero-Shot Learning (ZSL), Few-Shot Learning (FSL), Chain-of-Thought (CoT), Tree-of-Thought (ToT), and Guided Tree-of-Thought (GToT). ZSL uses minimal context for generalization, while FSL introduces examples to improve structure. For consistency, we use the same medium-sized class across experiments, ensuring token limit compliance for FSL examples. CoT enhances reasoning through step-by-step prompting, while ToT explores multiple reasoning paths before selecting test cases. GToT, our novel approach, simulates collaborative reasoning among three ”testers” who iteratively refine tests, combining CoT’s structured reasoning with ToT’s diversity to improve coverage, correctness, and variability. Figure 3 illustrates CoT and ToT prompts, while Figure 4 shows GToT’s collaborative structure. All prompts are available in our replication package.
@@ 3.5 The EvoSuite SBST tool
We employ its default settings, including the DynaMOSA algorithm (Panichella et al., 2017), which optimizes branch, method, and line coverage. To mitigate the randomness of genetic algorithms, we run EvoSuite 30 times per class, following (Tang et al., 2024). A three-minute time budget is allocated per test generation, aligning with prior studies (Fraser et al., 2015; Arcuri and Fraser, 2013; Shamshiri et al., 2015), which find this duration sufficient for effective test case generation. We deliberately disable EvoSuite’s post-generation processing, which reduces test smells, to ensure a fair comparison with unoptimized LLM-generated tests. This approach allows for an unbiased evaluation of the baseline capabilities of both methods, highlighting their inherent strengths and limitations.
@@ 3.6 Datasets
CMD includes modern, actively maintained open-source projects from the OceanBase Developer Center (ODC) and Conductor OSS, offering a complementary perspective to SF110 and Defects4J. While SF110 covers academic and industrial Java programs and Defects4J focuses on real-world bug-fixing scenarios, CMD introduces contemporary, evolving software. Despite its smaller scale, CMD is carefully curated to meet specific constraints for test generation evaluation. Its size is limited by three key factors. First, token constraints are considered, ensuring that each class can accommodate all prompt engineering techniques (ZSL, FSL, CoT, ToT, GToT) without exceeding token limits. Second, the dataset focuses on recent development, including only classes committed after May 2023 to reflect modern Java practices and prevent potential data leakage. Finally, CMD is designed to follow a buggy-fixed pairing approach, similar to Defects4J, where each class is selected with both a buggy and a fixed version, enabling structured and controlled evaluation of test generation performance. ODC (13 releases, 527 stars, 20 contributors) is a database development tool offering SQL development, risk management, and data security. Conductor OSS (11 releases, 15.5k stars, 53 contributors) is a microservices orchestration framework originally developed by Netflix. These modern, actively maintained projects provide a challenging testbed for LLM-generated tests, allowing us to assess generalization beyond the structured environments of SF110 and Defects4J. These projects are selected for their real-world relevance and evolving nature, reflecting modern Java development practices.
@@
SF110 Derived from DynaMOSA benchmark (Panichella et al., 2017), includes 346 Java classes from 117 projects. 74 182 170 138 120
Defects4J Contains 835 reproducible bugs across 17 open-source projects, focuses on detailed test suites to expose bugs. 16 477 254 177 130
Custom mini Dataset (CMD) Includes ODC, a collaborative project to enhance database functionalities, and Conductor OSS, an open-source project management tool aimed at efficiency and scalability. 2 31 16 9 5
@@
    Notice. Due to budgetary reasons, we only run GPT-4 and Mixtral-8x7B on the smallest dataset (CMD). Other models (GPT-3.5-Turbo and Mistral-7B) are run on all datasets.
@@ 4 Experimental Results
[Experiment Results]: Our study generated 216,300 files, with 190,619 (88.13%) containing source code. Table 3 summarizes MSR and CSR metrics across datasets and models, with detailed prompt-specific results in Tables 4 and 5.
@@
Table 3: MSR and CSR by LLM and Dataset.
Model Name # of Generated Files # of Files with code MSR CSR
SF110 Defects4J CMD SF110 Defects4J CMD SF110 Defects4J CMD SF110 Defects4J CMD
GPT 3.5-turbo 27 300 71 550 4 650 23 048 60 565 4 633 84.42% 84.65% 99.64% 78.10% 78.34% 90.79%
GPT 41 4 650 4 356 93.68% 83.59%
Mistral 7B 27 300 71 550 4 650 25 664 65 180 4 226 94.00% 91.10% 90.88% 79.97% 80.35% 83.61%
Mixtral 8x7B1 4 650 2 947 63.38% 53.57%
Total/Average (%) 54 600 143 100 18 600 48 712 125 745 16 162 89.21 87.88 86.90 79.04 79.35 77.89
@@
Finding 1: OpenAI models (GPT-3.5, GPT-4) demonstrate superior MSR and CSR, indicating their ability to generate structured, extractable test suites. Mistral 7B also performs well, whereas Mixtral 8x7B struggles, highlighting model-specific variations.
Impact of Prompting Techniques. GToT consistently achieved the highest MSR across all models and datasets(Table 4), reinforcing its effectiveness in enforcing structural adherence. For GPT-3.5-turbo, it significantly improved MSR, reaching 21.82% on Defects4J, 21.65% on SF110, and 21.94% on CMD. Mistral 7B also benefited from GToT, achieving 20.79% on Defects4J and 21.07% on SF110. Surprisingly, ZSL performed well in certain cases, with GPT-3.5-turbo reaching 21.98% MSR on CMD. In contrast, FSL consistently resulted in lower MSR scores, suggesting that additional examples do not necessarily enhance structural adherence. ToT outperformed CoT, improving MSR by up to 2.85% on SF110, reinforcing the advantage of stepwise reasoning in structuring outputs.
@@
Finding 2: GToT is the most effective prompting strategy for improving adherence to follow prompt instructions (MSR), while ZSL remains a strong baseline, especially for OpenAI models. FSL underperforms, challenging assumptions that more examples lead to better structural adherence.
GToT consistently achieved the highest CSR scores (Table 5), demonstrating its effectiveness in reducing extraneous content. GPT-3.5-turbo reached 21.10% on Defects4J, 19.65% on SF110, and 21.04% on CMD, while Mistral 7B followed a similar trend with 18.95% on CMD. ZSL remained a strong baseline, with GPT-3.5-turbo achieving 21.68% on CMD, indicating that direct instructions can sometimes outperform complex reasoning. FSL performed the worst, suggesting excessive prompt complexity introduces noise. Mixtral 8x7B struggled across all strategies, with its CSR dropping to 13.91% on CMD under GToT.
@@
GPT 3.5-turbo 18.5 14.13 15.05 17.21 21.10 17.69 14.37 15.0 16.37 19.65 21.68 20.75 20.07 16.37 21.04
GPT 4 21.15 21.06 19.03 16.19 14.55
Mistral 7B 17.87 17.85 18.21 17.76 16.52 18.08 17.99 17.82 17.38 13.81 18.95 18.67 18.74 18.5 17.13
Mixtral 8x7B 10.91 13.77 11.17 9.18 13.91
Finding 3: GToT improves extractability (CSR), reducing unnecessary content. However, ZSL remains competitive for GPT models, while FSL consistently underperforms. Mixtral 8x7B struggles significantly across all prompts.
To address LLMs’ tendency to mix test code with explanatory text, we introduced delimiters (###Test START### and ###Test END###) for structured output. Despite these clear boundaries, no LLM consistently follows the format, with MSR varying between 16% and 22% across models and datasets (Table 5). This inconsistency directly impacts CSR, which ranges from 9% to 22% (Table 5), highlighting persistent challenges in automated test code extraction.
Finding 4: Explicit delimiters improve extraction by structuring LLM outputs, but all LLMs exhibit inconsistencies in following formatting instructions, with some failing more frequently, making automated extraction challenging.
@@
[Experiment Goal]: This research question examines how prompt engineering influences test suite syntactical correctness and compilability. Valid Java grammar and compilation without modifications are essential for automated workflow integration. We aim to identify the most effective strategies for generating valid, executable test suites.
[Experiment Design]: We evaluate syntactical correctness using Javalang  [^6] https://github.com/c2nes/javalang parsing and assess compilability with JVM compilation. Tests are syntactically correct if parsing succeeds and compilable if compilation completes without errors. For non-compilable suites, we categorize compilation errors (e.g., ’Cannot Find Symbol’, ’Package Does Not Exist’) to identify hallucination patterns. We analyze how different prompting techniques affect these error patterns and assess their effectiveness in mitigating specific hallucinations, examining error distribution across strategies. This experiment spans SF110, Defects4J, and CMD datasets to evaluate how approaches influence correctness, compilability, and hallucination mitigation across diverse codebases.
@@
Dataset-Level Analysis. Table 3 presents a detailed breakdown of the LLM-generated test cases across SF110, Defects4J, and CMD. In total, our study generates 216,300 files, with 190,619 (88.13%) containing source code. Of these, 171,187 are syntactically correct, while 25,681 consist of textual outputs or structurally invalid test cases. These results highlight the challenge LLMs face in consistently generating well-formed unit tests, emphasizing the need for post-generation validation.
Finding 5: Approximately 21% of LLM-generated test suites that contain code fail syntactical correctness checks, requiring manual intervention before integration into real-world workflows.
LLM and Prompting Impact on Syntactical Correctness. Table 6 reveals significant variations in syntactical correctness across models and prompting techniques. GPT-3.5-turbo with GToT achieves the highest rates (23.65% on SF110, 24.54% on Defects4J, 21.06% on CMD), demonstrating structured reasoning’s benefits for code formation. GPT-4 reaches 23.00% on CMD using ZSL, indicating strong encoded syntactical patterns even with minimal prompting, while Mistral 7B performs best with FSL on SF110 (21.15%) and CoT on Defects4J (20.65%), suggesting different models benefit from different prompting approaches. Mixtral 8x7B consistently shows the lowest correctness rates across datasets.
Regarding prompting techniques, GToT consistently outperforms other strategies across models and datasets. Surprisingly, FSL underperforms despite providing examples, while ZSL performs well for GPT models, particularly GPT-3.5-turbo reaching 21.70% on CMD. These results highlight the crucial role of reasoning-based prompting in enhancing syntactical correctness, with different models showing varied responsiveness to specific techniques.
Finding 6: GPT models outperform others in syntactical correctness, with GPT-4 achieving strong results (23.00%) even with ZSL prompting, while Mixtral 8x7B struggles across datasets. GToT consistently improves correctness across all models, while FSL fails to enhance structure learning despite providing examples.
@@
Compilability Rates Across Datasets and Techniques. Table 7 reveals a significant gap between syntactical correctness and compilability. GPT-3.5-Turbo with GToT achieved the highest rates (9.67% on SF110, 6.76% on Defects4J), yet only 7.2% of generated tests and 12% of syntactically correct ones compiled. On CMD, GPT-3.5-Turbo with CoT surprisingly outperformed GPT-4 (5.92% vs. 2.65%), suggesting execution failures limit GPT-4’s syntax advantage. Mixtral 8x7B showed inconsistent performance (near-zero to 20.16% with GToT), indicating occasional success but poor overall reliability.

### v4

@@ 3.3 Large Language Models
Furthermore, all models are set to a standard temperature of 0.7. Our selection criteria prioritized models that excel in code generation and possess accessible APIs, making them suitable for both academic research and industrial applications (Wang et al., 2024a; Fan et al., 2023).
@@ 3.4 Prompt Design
This study evaluates five prompt techniques: Zero-Shot Learning (ZSL), Few-Shot Learning (FSL), Chain-of-Thought (CoT), Tree-of-Thought (ToT), and Guided Tree-of-Thought (GToT). ZSL uses minimal context for generalization, while FSL introduces examples to improve structure. For consistency, we use the same medium-sized class across experiments, ensuring token limit compliance for FSL examples. CoT enhances reasoning through step-by-step prompting, while ToT explores multiple reasoning paths before selecting test cases. GToT, our novel approach, simulates collaborative reasoning among three “testers” who iteratively refine tests, combining CoT’s structured reasoning with ToT’s diversity to improve coverage, correctness, and variability. Figure 3 illustrates CoT and ToT prompts, while Figure 4 shows GToT’s collaborative structure. All prompts are available in our replication package.
[…]
@@ 3.5 The EvoSuite SBST tool
For the SBST baseline, we selected EvoSuite (Fraser and Arcuri, 2011) as it is the most mature and empirically validated tool for automated test generation in Java. Recent editions of the SBFT competition confirm EvoSuite’s leadership over competitors such as Randoop (Pacheco and Ernst, 2007), Kex (Abdullin and Itsykson, 2022), and UTBot, particularly in terms of branch and line coverage (Jahangirova and Terragni, 2023). Large-scale empirical analyses (e.g. (Shamshiri et al., 2015)) also show that EvoSuite can detect real faults in Defects4J more effectively than random or commercial tools. Moreover, recent research on enhancing automated test generation (Ouédraogo et al., 2025) continues to adopt EvoSuite as the reference point when proposing extensions or hybrid approaches. These consistent results across competitions, benchmarks, and empirical studies justify the choice of EvoSuite as the most reliable SBST comparator for our study. We employ its default settings, including the DynaMOSA algorithm (Panichella et al., 2017), which optimizes branch, method, and line coverage. To mitigate the randomness of genetic algorithms, we run EvoSuite 30 times per class, following (Tang et al., 2024). A three-minute time budget is allocated per test generation, aligning with prior studies (Fraser et al., 2015; Arcuri and Fraser, 2013; Shamshiri et al., 2015), which find this duration sufficient for effective test case generation. We deliberately disable EvoSuite’s post-generation processing, which reduces test smells, to ensure a fair comparison with unoptimized LLM-generated tests. This approach allows for an unbiased evaluation of the baseline capabilities of both methods, highlighting their inherent strengths and limitations.
Unit of analysis.
Each EvoSuite run produces exactly one test suite, which corresponds to a single Java test file. Therefore, all counts and percentages reported in our study are based on the number of generated test files (one per suite). This ensures consistency with our treatment of LLM-generated outputs, where each model–prompt–iteration also produces one test suite (file).
@@ 3.6 Datasets
While SF110 covers academic and industrial Java programs and Defects4J focuses on real-world bug-fixing scenarios, both datasets were filtered with token constraints so that every retained class fits all prompt engineering techniques (ZSL, FSL, CoT, ToT, GToT). This reduced SF110 from 346 to 182 classes (74 projects) and Defects4J to 16 projects (477 classes). We built CMD to complement SF110 (academic/industrial programs) and Defects4J (bug-fixing scenarios) with contemporary, evolving software. It contains two actively maintained projects ODC (OceanBase Developer Center) and Conductor OSS, selected to assess generalization beyond legacy-style codebases. CMD is carefully curated under three constraints: (i) token limits so each class fits all prompt engineering techniques (ZSL, FSL, CoT, ToT, GToT), (ii) only classes committed after May 2023 to reflect modern Java practices and reduce potential training leakage, and (iii) a buggy–fixed pairing scheme (as in Defects4J) to enable controlled evaluation. ODC (13 releases, 527 stars, 20 contributors) is a database development tool providing SQL development, risk management, and data security; Conductor OSS (11 releases, 15.5k stars, 53 contributors) is a microservices orchestration framework originally developed at Netflix. Despite its smaller scale, CMD offers a modern, actively maintained testbed that complements SF110 and Defects4J.
@@
SF110 Derived from DynaMOSA benchmark (Panichella et al., 2017), originally 117 projects (346 classes). After applying token constraints to ensure all prompts fit within context windows, 74 projects (182 classes) were retained. 74 182 170 138 120
Defects4J Contains 835 reproducible bugs across 17 open-source projects. Similarly, token constraints were applied to guarantee that each selected class can accommodate all prompt engineering techniques without exceeding context limits, resulting in 16 projects (477 classes). 16 477 254 177 130
Custom mini Dataset (CMD) Includes ODC, a collaborative project to enhance database functionalities, and Conductor OSS, an open-source project management tool aimed at efficiency and scalability. To ensure fairness and prevent potential training leakage, CMD was curated under three constraints: (i) token limits, so that each class can be used across all prompt engineering techniques (ZSL, FSL, CoT, ToT, GToT), (ii) only classes committed after May 2023, and (iii) a buggy–fixed pairing approach similar to Defects4J. 2 31 16 9 5
@@
    Notice. Reported classes correspond to those retained after applying filtering constraints (token limits and, for CMD, post-2023 commits and buggy–fixed pairing). Due to budgetary reasons, we only run GPT-4 and Mixtral-8x7B on the smallest dataset (CMD). Other models (GPT-3.5-Turbo and Mistral-7B) are run on all datasets.
@@ 4 Experimental Results
4.1 [RQ0]: Ability of EvoSuite and LLMs to generate test code
[Experiment Goal]: This research question evaluates the fundamental ability of EvoSuite and LLMs to generate test suites, independently of syntactic correctness or compilability. The goal is to establish whether each paradigm can produce test code at scale for the target datasets, thereby clarifying their baseline feasibility as automated test generators.
[Experiment Design]: For each dataset and each class under test, we performed 30 independent generations. EvoSuite was applied on Defects4J and SF110 but not on CMD due to version compatibility issues with Java (Section 3.6). For LLMs, GPT-3.5-turbo and Mistral 7B were evaluated across all datasets, while GPT-4 and Mixtral 8×7B were restricted to CMD due to budget constraints (Section 3.6). The total number of expected test files was computed as the number of classes multiplied by 30 runs. For LLMs, since five prompting strategies were applied, the expected count was further multiplied by 5. Each generation, whether by EvoSuite or by an LLM, produces exactly one test suite, which corresponds to a single Java test file. Therefore, all counts and ratios in RQ0 are calculated on the basis of test files. The outcome metric is the ratio of generated test suites (i.e., outputs containing code) over the expected total, providing a direct measure of raw generation ability.
[Experiment Results]:
Table 3 summarizes the raw generation capacity of EvoSuite and the LLMs across datasets.
Table 3: Test generation capacity of EvoSuite and LLMs across datasets.
Dataset EvoSuite GPT-3.5 GPT-4 Mistral Mixtral
Defects4J 11100 / 14310 60565 / 71550 65180 / 71550
(77.6%) (84.7%) (91.1%)
SF110 5460 / 5460 23048 / 27300 25664 / 27300
(100%) (84.4%) (94.0%)
CMD 4633 / 4650 4356 / 4650 4226 / 4650 2947 / 4650
(99.6%) (93.7%) (90.9%) (63.4%)
  • 1
    Percentages refer to successful generations.
  • 2
    Mixtral = Mixtral 8×7B; Mistral = Mistral 7B; GPT-3.5 = GPT-3.5-turbo.
[…]
Relation to prior work.
Comparative studies with EvoSuite (e.g., Siddiq et al. (2024a); Tang et al. (2024); Yang et al. (2024)) typically report compilability or coverage after some code is returned, often on method-level targets or on filtered-class subsets (e.g., classes reduced to a single focal method with injected imports). In contrast, RQ0 isolates the raw question “does the tool emit an output file at all?” at the class level, where “per-class” means one generation per entire class under test, including all interdependent methods, with no import injection. Under this stricter setup across heterogeneous projects, we observe cases where EvoSuite produces no suite for some classes or projects, whereas LLMs usually emit some code output. This qualifies baseline feasibility and highlights the impact of granularity and dataset scope on raw generation.
Summary of RQ0: Even mature SBST tools like EvoSuite do not always return test suites for all classes, especially on heterogeneous datasets such as Defects4J. In contrast, LLMs almost always emit an output file, though at this stage we do not yet establish whether the content is valid test code. RQ0 therefore provides a baseline feasibility check: it clarifies whether tools emit outputs at all, before RQ1–RQ2 investigate format adherence, extractability, syntactic correctness, and compilability.
@@
These metrics are specific to LLMs, since unlike LLMs, EvoSuite directly produces JUnit test files without extraneous natural language or prompt-driven variability. As a result, MSR and CSR are not applicable to SBST tools and are only reported for LLM-generated outputs.
@@
[Experiment Results]:
Our study generated 216,300 files, with 190,619 (88.13%) containing source code.
Table 4 reports aggregated MSR and CSR scores by model and dataset, computed across all five prompting strategies (ZSL, FSL, CoT, ToT, GToT). This gives a global view of each model’s ability to produce extractable test code on average. The detailed breakdown by prompting technique is presented in Tables 5 and 6, allowing fine-grained analysis of how each prompt affects format compliance and extractability.
@@
Table 4: MSR and CSR by LLM and Dataset (aggregated over prompts)
. Model Name # of Generated Files # of Files with code MSR CSR SF110 Defects4J CMD SF110 Defects4J CMD SF110 Defects4J CMD SF110 Defects4J CMD GPT 3.5-turbo 27 300 71 550 4 650 23 048 60 565 4 633 84.42% 84.65% 99.64% 78.10% 78.34% 90.79% GPT 41 4 650 4 356 93.68% 83.59% Mistral 7B 27 300 71 550 4 650 25 664 65 180 4 226 94.00% 91.10% 90.88% 79.97% 80.35% 83.61% Mixtral 8x7B1 4 650 2 947 63.38% 53.57% Total/Average (%) 54 600 143 100 18 600 48 712 125 745 16 162 89.21 87.88 86.90 79.04 79.35 77.89
@@
Finding 2: OpenAI models (GPT-3.5, GPT-4) demonstrate superior MSR and CSR, indicating their ability to generate structured, extractable test suites. Mistral 7B also performs well, whereas Mixtral 8x7B struggles, highlighting model-specific variations.
Impact of Prompting Techniques. GToT consistently achieved the highest MSR across all models and datasets(Table 5), reinforcing its effectiveness in enforcing structural adherence.
For GPT-3.5-turbo, it reached 24.54% on Defects4J, 23.95% on SF110, and 20.01% on CMD. Mistral 7B also benefited from GToT, achieving 20.79% on Defects4J and 21.07% on SF110. GPT-4 reached 18.20% on CMD. ZSL showed competitive results in certain settings, with GPT-3.5-turbo reaching 21.51% on Defects4J and 20.05% on CMD. In contrast, FSL consistently produced lower MSR scores, suggesting that additional examples do not necessarily enhance structural adherence. ToT outperformed CoT across most models, improving MSR by up to 2.51% on Defects4J for GPT-3.5-turbo, highlighting the advantage of stepwise reasoning in structuring outputs.
@@
Finding 3: GToT is the most effective prompting strategy for improving adherence to follow prompt instructions (MSR), while ZSL remains a strong baseline, especially for OpenAI models. FSL underperforms, challenging assumptions that more examples lead to better structural adherence.
GToT often achieved the highest CSR scores across datasets (Table 6), particularly for GPT‑3.5‑turbo and GPT‑4, demonstrating its effectiveness in reducing extraneous content and structuring outputs. For example, GPT‑3.5‑turbo reached 24.54% on Defects4J and 23.65% on SF110 with GToT. However, CSR improvements are more sensitive to model–dataset interactions than MSR, and simpler prompting strategies can sometimes match or exceed GToT. On CMD, for instance, GPT‑3.5‑turbo achieved 21.70% with ZSL compared to 21.06% with GToT, and GPT‑4 also reached over 23% CSR with ZSL. For Mistral 7B, CoT slightly outperformed GToT on Defects4J (20.65% vs. 18.73%). FSL consistently underperformed across all settings, and Mixtral 8x7B remained the weakest overall despite peaking at 23.60% CSR on CMD under GToT.
@@
GPT 3.5-turbo 21.51 16.43 17.51 20.02 24.54 21.30 17.30 18.05 19.70 23.65 21.70 20.77 20.09 16.39 21.06
GPT 4 23.00 22.90 20.68 17.60 15.82
Mistral 7B 20.26 20.23 20.65 20.13 18.73 21.25 21.15 20.95 20.43 16.23 20.60 20.29 20.37 20.11 18.62
Mixtral 8x7B 18.51 23.36 18.95 15.58 23.60
Finding 4: GToT often yields the highest CSR scores, but improvements vary depending on the model and dataset. ZSL remains competitive for GPT models sometimes even outperforming GToT while FSL consistently underperforms. Mixtral 8x7B remains the least effective overall, despite isolated gains under GToT.
To address LLMs’ tendency to mix test code with explanatory text, we introduced delimiters (###Test START### and ###Test END###) to guide structured output. Despite these clear boundaries, no LLM consistently follows the expected format, with MSR varying between 16% and 25% across models and datasets (Table 6). This inconsistency directly impacts CSR, which ranges from 15% to 24% (Table 6), highlighting persistent challenges in reliably extracting complete and structured test code.
Finding 5: Explicit delimiters improve extraction by structuring LLM outputs, but all LLMs exhibit inconsistencies in following formatting instructions, with some failing more frequently, making automated extraction challenging.
Relation to prior work.
Existing studies on LLM-based test generation (e.g., Siddiq et al. (2024a); Tang et al. (2024); Yang et al. (2024); Yuan et al. (2024); Chen et al. (2024)) typically report results on syntactic correctness, compilability, coverage, or fault detection once code has been extracted, without explicitly assessing the reliability of that extraction step. Their evaluation pipelines therefore proceed directly from raw responses to downstream metrics, leaving output format and extractability unexamined. In this study we adapt the MSR/CSR framework of Macedo et al. (2024), originally proposed for code translation, to JUnit test generation, making format and extractability an explicit precondition before syntax and compilation analyses (RQ2).
Summary of RQ1: Prompt engineering strongly influences extractability of LLM-generated test suites. OpenAI models (GPT-3.5-turbo, GPT-4) achieve the highest MSR and CSR, with Mistral 7B competitive and Mixtral 8×7B lagging. GToT consistently yields the highest MSR across all models and datasets, enforcing structural adherence. For CSR, GToT often attains the best results (particularly for GPT models), though simpler prompts like ZSL or CoT occasionally match it depending on model–dataset pairing. FSL underperforms across all settings. These findings extend Macedo et al. (2024) by showing that controlling prompt structure remains crucial in test generation: explicit reasoning and delimiter guidance substantially enhance format compliance, yet no model follows delimiters reliably. Extracting usable test code from mixed natural-language outputs remains a key challenge for practical adoption.
@@
[Experiment Goal]: This research question examines how prompt engineering influences test suite syntactical correctness and compilability in LLM-generated tests. Valid Java grammar and compilation without modifications are essential for automated workflow integration. We aim to identify the most effective prompting strategies for generating valid, executable test suites. To provide a reference point, we also report EvoSuite’s syntactic correctness and compilability, as it represents a state-of-the-art SBST baseline. This contrast highlights differences between search-based and LLM-based approaches in producing directly usable test code.
[Experiment Design]: We evaluate syntactical correctness using Javalang [^13] https://github.com/c2nes/javalang parsing and assess compilability with JVM compilation. A test suite is considered syntactically correct if parsing succeeds, and compilable if compilation completes without errors. For non-compilable suites, we categorize compilation errors (e.g., Cannot Find Symbol, Package Does Not Exist) to identify hallucination patterns. We then analyze how different prompting techniques affect these error distributions and assess their effectiveness in mitigating specific hallucinations. EvoSuite results are included to indicate the upper bound expected from an SBST tool, allowing us to contextualize LLM outputs against a mature baseline. Experiments span SF110, Defects4J, and CMD datasets to evaluate how both paradigms handle correctness, compilability, and hallucination mitigation across diverse codebases.
Each run, whether by an LLM or by EvoSuite, produces exactly one test suite, which corresponds to a single Java test file. Therefore, all counts and percentages in our study are calculated on the basis of test files, ensuring a consistent comparison between LLM-generated and SBST-generated outputs.
For the calculation of syntactic correctness and compilability percentages, we proceed as follows. Suppose $A$ is a given model and $B$ a given dataset. Let $X$ be the total number of syntactically correct test suites produced by $(A,B)$ across all prompting strategies. For a specific prompting strategy $p$, we denote by $Y_{p}$ the number of syntactically correct test suites it generates, and by $Z_{p}$ the number of those that also compile successfully. The syntactic correctness percentage for $(A,B,p)$ is then computed as $Y_{p}/X\times 100$, while the compilability percentage is computed as $Z_{p}/X\times 100$. For EvoSuite, which does not involve prompting strategies, syntactic correctness is measured as the fraction of syntactically correct test suites among all generated ones, and compilability is measured as the fraction of compilable test suites among those that are syntactically correct.
@@
Across all LLM runs, we produced 216,300 outputs (Table 3,Table 4). Of these, 190,619 (88.13%) contained extractable test code, while the remaining 25,681 (11.87%) were plain text or structurally invalid (no extractable code). Among the code-containing outputs, 171,187 (89.81%) were syntactically correct, while 19,432 (10.19%) exhibited syntax errors. For EvoSuite, all generated test files (16,560 across Defects4J and SF110) were syntactically correct. These results highlight the challenge LLMs face in consistently generating well-formed unit tests, emphasizing the need for post-generation validation.
Finding 6: Across datasets, 89.81% of LLM-generated outputs containing code are syntactically correct, while 10.19% include syntax errors, requiring manual intervention before integration into real-world workflows. In contrast, EvoSuite achieves 100% syntactic correctness on all its generated outputs.
To provide full transparency, Table 7 reports the raw counts of syntactically correct (S) and compilable (C) test suites for each model, dataset, and prompting strategy. These raw values form the basis for the normalized percentages of syntactic correctness and compilability later reported in Tables 8 and 9.
Table 7: Syntactic correctness (S) and compilability (C) by model, dataset, and prompt.
Model Dataset ZSL (S/C) FSL (S/C) CoT (S/C) ToT (S/C) GToT (S/C)
GPT-3.5-turbo Defects4J 12056/1656 9207/1244 9812/1808 11220/1268 13756/3789
SF110 4541/1267 3689/897 3849/1164 4200/870 5042/2061
CMD 916/114 877/103 848/250 692/113 889/146
GPT-4 CMD 894/69 890/86 804/103 684/74 615/93
Mistral 7B Defects4J 11649/258 11633/247 11871/153 11575/252 10766/629
SF110 4639/238 4617/227 4574/199 4460/379 3543/735
CMD 792/0 789/1 801/0 782/1 724/12
[…]
EvoSuite Defects4J 11100/10890
SF110 5460/5460
S = number of tests that parse successfully (syntactically correct); C = number of tests that compile without errors. Each cell reports S/C.
EvoSuite Compilation Outcomes. While EvoSuite achieved full syntactic correctness on all the test files it generated, a closer inspection reveals that not all of these tests compiled successfully. In particular, we identified 210 non-compilable cases across Defects4J: two in JacksonXml, two in Math, and three in JacksonDatabind. This represents a small fraction of the total (210 out of 11,100 syntactically correct suites, i.e., 1.9%), yet it highlights that even mature SBST tools can occasionally produce outputs that fail at compilation. On SF110, by contrast, EvoSuite compiled all 5,460 generated test files without error.
Finding 7: Even EvoSuite, the SBST baseline, does not guarantee full compilability: although all its generated tests were syntactically correct, 210 cases in Defects4J ($\approx 1.9\%$) failed to compile, while SF110 compiled without errors. This shows that compilation failures are not unique to LLMs but can also occur in mature automated test generation tools.
LLM and Prompting Impact on Syntactical Correctness. Table 8 reveals significant variations in syntactical correctness across models and prompting techniques. GPT-3.5-turbo with GToT achieves the highest correctness on Defects4J (24.54%) and SF110 (23.65%), showing the benefit of structured reasoning for syntax. On CMD, however, it performs slightly lower than ZSL (21.06% vs. 21.70%). GPT-4 also reaches high scores on CMD with ZSL (23.00%). For Mistral 7B, FSL performs best on SF110 (21.15%), while CoT is highest on Defects4J (20.65%). These results suggest that prompting effectiveness for syntactical correctness varies across models and datasets, with no single strategy universally superior. FSL remains inconsistent overall, and Mixtral 8x7B shows the weakest performance despite isolated peaks (e.g., 23.60% with GToT on CMD).
Finding 8: GPT models outperform other LLMs in syntactic correctness, with GPT-4 achieving the highest correctness rate (23%) among all prompt configurations. Prompting effectiveness, however, is dataset-dependent: GToT yields top results on Defects4J and SF110, yet ZSL slightly surpasses it on CMD. No single prompting strategy proves universally superior, while Mixtral 8x7B remains the weakest performer overall despite occasional peaks (e.g., 23.6% with GToT on CMD).
@@
Compilability Rates Across Datasets and Prompting Techniques. Table 9 reveals a significant gap between syntactical correctness and compilability. GPT-3.5-Turbo with GToT achieved the highest rates (9.67% on SF110, 6.76% on Defects4J), yet only 7.2% of generated tests and 12% of syntactically correct ones compiled. On CMD, GPT-3.5-Turbo with CoT surprisingly outperformed GPT-4 (5.92% vs. 2.65%), suggesting execution failures limit GPT-4’s syntax advantage. Mixtral 8x7B showed inconsistent performance (near-zero to 20.16% with GToT), indicating occasional success but poor overall reliability.